OpenStudy (anonymous):

Cards numbered 1, 1, 2, 3, 4, 4, 4, 5, 6, 6, 7, 8, 8, 8, and 9 are placed in a bag. If a card is drawn from the bag and then replaced 135 times, about how many times would a 6 be drawn? A.2 B.9 C.15 D.18

OpenStudy (anonymous):

The probability of picking a 6 is 2/15 correct?

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

no that is only for 1 time

OpenStudy (anonymous):

If you are doing it 135 times you should multiply 2/15 by 135

OpenStudy (anonymous):

b

OpenStudy (anonymous):

As the number of draws increases arbitrarily, you should expect the proportion of 6s drawn to be the same as the proportion of 6s in the deck. In this case, you have two 6s in a deck of fifteen cards, which means for any given draw, you have a \(\dfrac{2}{15}\) chance of drawing a 6. (Since the card is replaced, you have an equal probability of drawing a 6 the second time around, and so on for all remaining draws.) Now, if you draw \(n\) times, you should expect to get \(\dfrac{2}{15}n\) 6s (as \(n\to\infty\)).

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

2/15*135= 270/15= 18...